<html>
  <head>

    <meta http-equiv="content-type" content="text/html; charset=ISO-8859-1">
  </head>
  <body>
    <meta http-equiv="content-type" content="text/html;
      charset=ISO-8859-1">
    BTeamers,<br>
    <br>
    Here is the agenda for tomorrow's 11:00 BTeam meeting. We had almost
    no studies last week, so this<br>
    week's study period is a repeat of last week's schedule. We'll also
    review the latest round of machine<br>
    tuning ongoing today, and have a peek at our readiness for 5-pass
    spinup and some comments from<br>
    Tief on his elog entry about measuring bunch length with the 1A sync
    light monitor. We'll see you<br>
    at the BTeam meeting tomorrow! Cheers!<br>
    <br>
    -Todd<br>
    <br>
    <h1 class="title">Jan 24 2012 BTeam Meeting Agenda</h1>
    <div class="outline-2">
      <h2 id="sec-2">1 Optics on Call Report/Schedule</h2>
      <ul>
        <li>Rotates at BTeam (10:00 Tuesdays)<br>
        </li>
        <li>Medium-term machine development<br>
          <ul>
            <li><span class="done">DONE</span> Write 5-pass spin-up
              ATLis (Jay, <a
                href="https://cebaf.jlab.org/atlis/task/11894">a11894</a>)<br>
            </li>
          </ul>
        </li>
        <li>17 Jan-24 Jan: Vasily/Yves (Outgoing)<br>
          <ul>
            <li>SL has still been running few degrees off crest to
              reduce Compton backgrounds (Jay)<br>
            </li>
            <li>Hall B sensitivity to energy drift of some concern
              (Tief)<br>
            </li>
            <li><span class="done">DONE</span> Mon 23 Jan Add M56 optics
              in injector (Yves/Jay, <a
                href="https://cebaf.jlab.org/atlis/task/11924">a11924</a>)<br>
            </li>
            <li><span class="todo">TODO</span> Restore SL MOMOD after SL
              on crest, good hall C Compton rates confirmed (Tief/ops)<br>
            </li>
          </ul>
        </li>
        <li>24 Jan-31 Jan: Ed/Alex (Oncoming)<br>
        </li>
        <li>31 Jan-07 Feb: Yuhong/Geoff<br>
          <ul>
            <li>5-pass spin-up schedule? Feb 1?<br>
            </li>
          </ul>
        </li>
        <li>07 Feb-14 Feb: Balsa/Todd<br>
        </li>
        <li>14 Feb-21 Feb: Fanglei/Mike Spata<br>
        </li>
        <li>21 Feb-28 Feb: Todd/Jay<br>
        </li>
      </ul>
    </div>
    <div class="outline-2">
      <h2 id="sec-3">2 Measuring Bunch Length using 1A Sync Light
        Monitor (Tief)</h2>
      <ul>
        <li>New BTeam elog entry on this procedure: <a
            href="https://cebaf.jlab.org/elog/entry/1661285">e1661285</a><br>
        </li>
      </ul>
    </div>
    <div class="outline-2">
      <h2 id="sec-4">3 C100 Status</h2>
      <ul>
        <li>26/52/79 Hz signal on 1/4 cryo drive gradient/phase
          (Suleiman/Spata)<br>
        </li>
        <li>Beam studies rescheduled for Wed 25 Feb<br>
        </li>
      </ul>
    </div>
    <div class="outline-2">
      <h2 id="sec-5">4 Beam Studies Report 18 Jan 2012</h2>
      <ul>
        <li>Studies scheduled 08:00-20:00 (on time)<br>
          <ul>
            <li> <a href="https://cebaf.jlab.org/atlis/task/11858">a11858</a>
              0.0 C100: Search for 79 Hz (Parisitic, Suleiman/Spata;
              strong 52 Hz on 1/4 cryo drive gradient/phase) </li>
            <li> No other studies accomplished due to box supply
              failures, repair </li>
          </ul>
        </li>
      </ul>
    </div>
    <div class="outline-2">
      <h2 id="sec-6">5 Beam Studies Scheduled 25 Jan 2012</h2>
      <ul>
        <li><span class="todo">TODO</span> Submit ATLis items for
          desired beam studies for next week (25 Jan)<br>
        </li>
        <li>Same studies as last week based on lost time to box supply
          failures<br>
        </li>
        <li>Studies scheduled 08:30-20:30<br>
          <ul>
            <li> <b>[<span style="visibility:hidden;">X</span>]</b> ——
              1.0 Optics: Fopts/allsaves </li>
            <li> <b>[<span style="visibility:hidden;">X</span>]</b> <a
                href="https://cebaf.jlab.org/atlis/task/11882">a11882</a>
              8.0 C100: Stability testing (Hovater, Spata, etc, full
              gradient on both C100s, 10uA) </li>
            <li> <b>[<span style="visibility:hidden;">X</span>]</b> <a
                href="https://cebaf.jlab.org/atlis/task/11865">a11865</a>
              1.0 Modeling: RayTrace validation (Bodenstein, Tief) </li>
            <li> <b>[<span style="visibility:hidden;">X</span>]</b> <a
                href="https://cebaf.jlab.org/atlis/task/11885">a11885</a>
              0.0 PePPo: Modeling (Grames, Contingency) </li>
            <li> <b>[<span style="visibility:hidden;">X</span>]</b> <a
                href="https://cebaf.jlab.org/atlis/task/11849">a11849</a>
              0.0 Mott: Mott DAQ Test (Suleiman, Contingency) </li>
            <li> <b>[<span style="visibility:hidden;">X</span>]</b> ——
              2.0 Optics: Restoration </li>
          </ul>
        </li>
        <li><span class="todo">TODO</span> Optics should write notes,
          observations of fopt/etc to Bteam elog<br>
          <ul>
            <li>Include pathlength check (pathlength was good Mon day)<br>
            </li>
          </ul>
        </li>
      </ul>
    </div>
    <div class="outline-2">
      <h2 id="sec-7">6 Next week</h2>
      <ul>
        <li>Alex Bogacz has graciously agreed to chair while Todd is on
          travel<br>
        </li>
        <li>Hall B to 5th pass ~ Feb 1<br>
        </li>
      </ul>
    </div>
    <div class="outline-2">
      <h2 id="sec-8">7 From last week: Evaluate possibilities of Krest
        during beam restoration (Okay)</h2>
      <ul>
        <li>Takes more like 4h to run through both linacs to get decent
          results (Terry)<br>
        </li>
        <li>But Hall C doesn't get good data while Krest is running
          either (but Hall B can)<br>
        </li>
        <li>Perhaps part of 2h can be a manual touch-up of cresting (15
          min total, 2 operators)<br>
          <ul>
            <li><span class="todo">TODO</span> Todd will discuss with
              Noel to see if manual touch-up is acceptable weekly retune<br>
            </li>
          </ul>
        </li>
      </ul>
    </div>
    <div id="postamble"><br>
    </div>
    <meta http-equiv="content-type" content="text/html;
      charset=ISO-8859-1">
    <title>BTeam-2012-01-24</title>
    <meta http-equiv="Content-Type" content="text/html;
      charset=ISO-8859-1">
    <meta name="generator" content="Org-mode">
    <meta name="generated" content="2012/01/23 11:26:18">
    <meta name="author" content="Todd Satogata">
    <style type="text/css">
  html {
        font-family: Times, serif;
        font-size: 12pt;
  }
  .title { text-align: center; }
  .todo  { color: red; }
  .done { color: green; }
  .timestamp { color: grey }
  .timestamp-kwd { color: CadetBlue }
  .tag { background-color:lightblue; font-weight:normal }
  .target { background-color: lavender; }
  pre {
        border: 1pt solid #AEBDCC;
        background-color: #F3F5F7;
        padding: 5pt;
        font-family: courier, monospace;
  }
  table { border-collapse: collapse; }
  td, th {
        vertical-align: top;
        <!--border: 1pt solid #ADB9CC;-->
  }
</style>
  </body>
</html>